# 100,000 Crab giant pulses at 215 MHz detected with an SKA-Low prototype station

### v1

@@ Abstract
We report detection and analysis of the largest ever low-frequency sample of Crab giant pulses (GPs) detected in frequency band 200 – 231.25 MHz. In total about 95000 GPs were detected, which, to our knowledge is the largest low-frequency sample of Crab GPs presented in the literature. The observations were performed between 2024-12-14 and 2025-03-31 with the Engineering Development Array 2, a prototype station of the low-frequency Square Kilometre Array telescope. The fluence distribution of GPs in the entire sample is very well characterised with a single power law N(F) Fα, where for all GPs, and and for GPs at the phases of the main pulse and interpulse respectively. We do not observe flattening of the fluence distribution at the higher fluences. Although, the index of the power law fluence distribution remained approximately constant over the observing period, the normalisation of the distribution was strongly anti-correlated (coefficient ) with the scatter broadening time. These variations were observed on the timescale of weeks, pointing to refractive scintillation as the root cause. As a result, the measured fluence distribution was augmented for lower ( 2 ms) and diminished for higher ( 5 ms) scatter broadening time causing the GP detection rate to vary between 3000 and 100 per hour respectively (the correlation coefficient -0.9). Furthermore, for the first time at low-frequencies we observe a strong positive correlation (correlation coefficient 0.6) between the scatter broadening time () and dispersion measure. Our modelling favours the screen size pc with mean electron density e-cm-3 located within 100 pc from the pulsar (Crab Nebula or Perseus arm of the Milky Way galaxy). The observed frequency scaling of the scattering broadening time (where ) is in agreement with the previous measurements. The observed maximum spectral luminosities erg/Hz/s approach those of the weakest pulses from some repeating Fast Radio Bursts (FRBs). We also find evidence of deviation of GP time separation from a purely random (Poisson) process. Overall, our results agree with the current views that GPs from extra-galactic Crab-like pulsars can be responsible for some very weak repeating FRBs, but cannot explain the entire FRB population. Finally, these results demonstrate an enormous transient science potential of individual SKA-Low stations, which can be unlocked by milli-second all-sky imaging.
@@ 1 Introduction
Pulsar B0531+21, aka Crab, is one of the first discovered pulsars (1968Sci...162.1481S). It is also one of the most interesting and longest studied pulsars capturing the imagination of the general public due to the magnificent Crab Nebula. It is a truly fascinating object and one of the “living proofs” that supernova (SN) explosions can lead to formation of a neutron star as the Crab pulsar and the surrounding nebula were formed in the core-collapse SN 1054 observed by medieval astronomers. As such, the Crab pulsar is considered to be a young pulsar with a period ms. It is located at the distance of approximately 1.9 kpc (2023ApJ...952..161L) and moving with a velocity of about 120 km s-1 (2008ApJ...677.1201K). The average profile of the Crab pulsar changes significantly as a function of frequency, and exhibits the main pulse (MP) with a precursor, and an interpulse (IP) at about 145°offset (about 13 ms) from the MP.
@@
The Crab pulsar is one of a very few pulsars emitting the so-called giant pulses (GPs) - extremely narrow (ns) pulses with brightness temperatures reaching K (2003Natur.422..141H; 2004ApJ...612..375C). Crab GPs were discovered shortly after the discovery of the pulsar (1970Natur.226..529H; 1970Natur.226...69S). The physical mechanisms behind the emission of GPs remain unexplained (see review by 2016JPlPh..82c6302E). However, it is well established that the GP energies follow a power law distribution (1972ApJ...175L..89A; 2004ApJ...612..375C; 2008ApJ...676.1200B; 2015PASA...32....6O; 2017ApJ...851...20M; 2019MNRAS.490L..12B) in contrast to the exponential or log-normal distribution of normal pulses (2012MNRAS.423.1351B), which indicates different emission mechanisms.
@@
After the discovery of fast radio bursts (FRBs; 2007Sci...318..777L), GPs from extra-galactic pulsars were hypothesised to be responsible for at least some types (i.e. repeating) of FRBs (2016MNRAS.457..232C), and Crab GPs have been extensively studied in the context of FRBs (e.g. 2022NatAs...6..393N; 2019ApJ...876L..23H) as well as for testing this hypothesis (e.g. 2024ApJ...973...87D; 2019MNRAS.490L..12B; 2017ApJ...851...20M). Based on the energy distribution of Crab GPs and other characteristics (e.g. time between pulses), these studies have disfavoured the theory that all FRBs are caused by GPs from extra-galactic pulsars, and it is more likely that majority of FRBs are caused by objects like magnetars (2019MNRAS.485.4091M; 2021A&A...651A..63G). However, some weaker FRBs, such as FRB 20200120E, could be caused by extra-galactic Crab-like pulsars (see 2022NatAs...6..393N). For the most recent review of the FRB field see 2022A&ARv..30....2P, while FRB models are collated at https://frbtheorycat.org/.
@@
While the discovery and early Crab GP studies (e.g. 1995ApJ...453..433L) were performed at frequencies below 1 GHz, the majority of later investigations were conducted at frequencies above 1 GHz (e.g. 2024ApJ...973...87D; 2019MNRAS.490L..12B; 2019MNRAS.483.1224D; 2008ApJ...676.1200B). On the high end of electromagnetic spectrum GP studies extend to -rays; see 2021Univ....7..448A for the most recent review, and 2021Sci...373..425L; 2023ApJ...956...80A for recent results at PeV energies.
@@
The advent of new (LOFAR (2013A&A...556A...2V), MWA (2013PASA...30....7T; 2018PASA...35...33W), LWA (2013ITAP...61.2540E) etc.) and upcoming instruments, like low-frequency Square Kilometre Array telescope (SKA-Low; 2009IEEEP..97.1482D)111https://www.skatelescope.org/ with large fields-of-view (FoVs), has begun to unlock the wealth of information available at these frequencies, and reinvigorated pulsar studies at frequencies below 400 MHz. Several recent GP studies were performed down to 110 MHz (e.g. 2012A&A...538A...7K; 2017ApJ...851...20M; 2015ApJ...809...51O; 2012ApJ...760...64M; 2007ApJ...665..618B). These studies addressed various aspects of Crab GPs ranging from monitoring plasma in the surrounding environment and along the line of sight (LoS) to the pulsar, through variations in dispersion measure (DM) and scatter broadening time to statistical properties of GPs, their emission mechanisms and similarities to FRBs.
@@
This paper presents results from one such instrument, the Engineering Development Array version 2 (EDA2; 2022JATIS...8a1010W) a prototype station of the SKA-Low. Despite modest sensitivity we were able to record the largest ever sample (95000) of Crab giant pulses observed at SKA-Low frequencies, superseding even the largest ever sample of 30000 Crab GPs recorded by 1995ApJ...453..433L. The EDA2 was originally deployed as a risk mitigation and reference system to the official SKA-Low prototype station the Aperture Array Verification System 2 station (AAVS2; 2022JATIS...8a1014M). Real-time station beamforming and pulsar observing mode were commissioned at an early stage and already resulted in early SKA-Low pulsar science results (2022PASA...39...42L; 2024PASA...41...80L). Both stations have also been used for all-sky transient searches (2021PASA...38...23S).
@@
Recently, a new pipeline, based on a code developed for the MWA by 2024PASA...41...11S, has been implemented and deployed for non-targeted (“blind”) searches and monitoring of repeating FRBs, magnetars, RRATs and pulsars with a single station beam of EDA2. In order to test and verify the pipeline, multiple observations of Crab pulsar were performed, and led to detection of thousands of GPs. These detections have confirmed that the station and the pipeline are capable of detecting highly dispersed pulses, and can be successful in searching for and/or monitoring FRBs, magnetars, RRATs or other fast transients. Especially, once the full FoV of the stations is unlocked by multi-beaming or high-time resolution all-sky imaging as proposed by 2022aapr.confE...1S.
@@ 2 EDA2 telescope and observations
The data for this study were recorded between 2024-12-14 and 2025-03-31 with the EDA2 station. The details of EDA2 can be found in the description paper (2022JATIS...8a1010W), and only a brief summary relevant to this project is provided here. The EDA2 has been a risk mitigation and a reference to the AAVS2 station, with the only major difference being the antenna design. The EDA2 comprises 256 MWA bow-tie dipoles, and was used as a reference to AAVS2 station (already de-commissioned) composed of 256 SKALA 4.1 antennas. Both instruments were located close to each other (some 500 m apart) at the SKA-Low site, Inyarrimanha Ilgari Bundara observatory222https://www.csiro.au/en/about/facilities-collections/mro. Analogue signals from individual antennas are digitised and the entire bandwidth (400 MHz) channelised into 512 coarse channels in the backend Tile Processing Units (TPMs; 2017JAI.....641014N). One of the main tasks of TPMs is to form real-time station beams in 1.08 usec time resolution over the entire SKA-Low observing bandwidth (50 – 350 MHz).
@@ 3 Data processing pipeline
The recorded DADA files have been processed with eda2frb pipeline333https://github.com/marcinsokolowski/eda2frb developed for this project (Fig. 1). The pipeline has been based on the FRB search pipeline applied to MWA incoherent beam (2024PASA...41...11S)444athttps://github.com/marcinsokolowski/mwafrb. It has been executed on a server computer installed at the observatory and consists of the following main steps. First, complex voltages in each coarse channel DADA file are fine channelised into 64 fine channels (approximately 14.5 kHz each) by a custom-developed spectrometer using FFTW (CPU) or cuFFT (GPU) libraries (the choice is optional). The resulting spectra (in 69.12 usec time resolution) are averaged to 0.96768 ms time resolution, and saved to separate filterbank files. Then, 40 filterbank files from each coarse channel are merged into a single filterbank covering the entire observing bandwidth (31.25 MHz). These steps are represented by blue boxes in Figure 1.
The single pulse search is performed on the resulting wideband filterbank file using both PRESTO (2011ascl.soft07017R) and FREDDA (2019ascl.soft06003B) software packages. One of the PRESTO executions was performed at Crab pulsar’s DM=56.72 pc/cm3 as measured during this period by the Jodrell Bank observatory555https://www.jb.man.ac.uk/pulsar/crab/crab2.txt. Selected FREDDA candidates were visually inspected, and an example of a GP with the highest signal-to-noise (SNR) ratio of 354 recorded in this campaign is shown in Figure 2. This is the brightest pulse in the sample with flux density of 16.6 kJy and fluence 76 Jy ms. We note that the pulse broadening effects were not removed in our analysis (e.g. via de-convolution, 2003ApJ...584..782B), which may lead to a factor of 2 underestimation of the intrinsic peak flux density (see discussion in 2007ApJ...665..618B).
@@ 4.1 Post-processing of PRESTO output
Merging candidates - a single GP may result in multiple candidates with different arrival times, DMs and SNRs. Therefore, in order to avoid double-counting, the candidates were merged into a single candidate with maximum SNR within a 45 ms time window as highly scattering-broadened pulses can extend up to this duration. In principle, we could have missed two GPs within one pulsar period (33 ms), but the results of merging were visually inspected and such cases have never been found. In fact, so far, the so-called double (or “twin”) GPs, from MP and IP within the same pulsar period, have only been reported in L-band (2010A&A...515A..36K). Therefore, we also performed a dedicated search for such events in every recorded de-dispersed time series, and did not find any instance of two pulses separated by less than a pulsar period (33 ms). Hence, we are confident that the merging procedure does not affect the statistical analysis of fluence and flux density distributions, or any other parts of the presented analysis.

### v2

@@ Abstract
We report detection and analysis of the largest ever low-frequency sample of Crab giant pulses (GPs) detected in frequency band 200 – 231.25 MHz. In total about 95000 GPs were detected, which, to our knowledge is the largest low-frequency sample of Crab GPs presented in the literature. The observations were performed between 2024-12-14 and 2025-03-31 with the Engineering Development Array 2, a prototype station of the low-frequency Square Kilometre Array telescope. The fluence distribution of GPs in the entire sample is very well characterised with a single power law N(F) Fα, where for all GPs, and and for GPs at the phases of the main pulse and low-frequency interpulse respectively. We do not observe flattening of the fluence distribution at the higher fluences. Although, the index of the power law fluence distribution remained approximately constant over the observing period, the normalisation of the distribution was strongly anti-correlated (coefficient ) with the scatter broadening time. The timescale ( weeks) of these variations indicates that intrinsic GP emission was modulated by the refractive scintillation as the signals propagated through the Crab Nebula and ISM. As a result, the measured fluence distribution was augmented for lower ( 2 ms) and diminished for higher ( 5 ms) scatter broadening time causing the GP detection rate to vary between 3000 and 100 per hour respectively (the correlation coefficient -0.9). Furthermore, for the first time at low-frequencies we observe indications of positive correlation (correlation coefficient 0.7) between the scatter broadening time () and dispersion measure. Our modelling favours the screen size pc with mean electron density e-cm-3 located within 100 pc from the pulsar (Crab Nebula or Perseus arm of the Milky Way galaxy). The observed frequency scaling of the scattering broadening time (where ) is in agreement with the previous measurements. The observed maximum spectral luminosities erg/Hz/s approach those of the weakest pulses from some repeating Fast Radio Bursts (FRBs). However, the distribution of pulse arrival times is consistent with a purely random Poisson process, and we do not find evidence of clustering. Overall, our results agree with the current views that GPs from extra-galactic Crab-like pulsars can be responsible for some very weak repeating FRBs, but cannot explain the entire FRB population. Finally, these results demonstrate an enormous transient science potential of individual SKA-Low stations, which can be unlocked by milli-second all-sky imaging.
@@ 1 Introduction
Pulsar B0531+21, aka Crab, is one of the first discovered pulsars 1968Sci…162.1481S (85). It is also one of the most interesting and longest studied pulsars capturing the imagination of the general public due to the magnificent Crab Nebula. It is a truly fascinating object and one of the “living proofs” that supernova (SN) explosions can lead to formation of a neutron star as the Crab pulsar and the surrounding nebula were formed in the core-collapse SN 1054 observed by medieval astronomers. As such, the Crab pulsar is considered to be a young pulsar with a period ms. It is located at the distance of approximately 1.9 kpc 2023ApJ…952..161L (51) and moving with a velocity of about 120 km s-1 2008ApJ…677.1201K (38, 51).
@@
The Crab pulsar is one of a very few pulsars emitting the so-called giant pulses (GPs). These are very short (s) bursts often exhibiting nanosecond substructure consisting of extremely narrow nano-shots (e.g., see 2007ApJ…670..693H (32) or 2010A&A…524A..60J (37)), which implies brightness temperatures reaching K 2003Natur.422..141H (33, 15). Crab GPs were discovered shortly after the discovery of the pulsar 1970Natur.226..529H (34, 84). Their energies follow a power law distribution 1972ApJ…175L..89A (3, 15, 10, 67, 63, 5) in contrast to the exponential or log-normal distribution of normal pulses observed in other pulsars 2012MNRAS.423.1351B (13), which indicates different emission mechanisms.
Interestingly, Crab average pulse profiles are considered to be an average of a large ensemble of GPs (e.g. 2021ApJ…920…38B (12, 60)). Hence, even though the physical mechanisms behind the GP emission remain unexplained (see review by 2016JPlPh..82c6302E (24)), they make the Crab pulsar a unique laboratory to study pulsar emission mechanisms (for example, see discussions in 2007ApJ…670..693H (32, 37, 12)).
Crab’s average pulse profile changes significantly as a function of frequency, and 8 components have been reported in literature (e.g. see 2023ApJ…959..111L (50) for a detailed summary). At frequencies below 2 GHz and at higher photon energies (from near infrared up to hard -rays) the average pulse profile consists of the main pulse (MP) with a precursor and a low-frequency interpulse (IP) at a phase offset of about 145° (about 13 ms) from the MP. At these frequencies the average pulse profiles are consistent with the standard model that MPs and IPs originate from physically very similar regions on the opposite sides of the neutron star (NS). However, this picture becomes much more complicated at frequencies 2 – 10 GHz. Firstly, at around 2.7 GHz the IP component disappears, while at higher frequencies (4 GHz) the IP re-appears, but shifted in phase by about 20° from the low-frequency IP. Therefore, it is believed to be a different component, and referred to as high-frequency IP. Furthermore, two additional high-frequency components (HFC1 and HFC2) show up at pulse phases following the IP (see Figure 1 in 2007ApJ…670..693H (32)). Finally, 2007ApJ…670..693H (32) discovered that properties (temporal and spectral structures, DM and polarisation) of high-frequency GPs occurring at phases of MP and high-frequency IP differ significantly. This implies that the physical mechanisms and conditions in the emission region of the high-frequency IP are very different to those of the MP - in contrast to standard pulsar model where their are expected to be nearly the same just on the opposite sides of the NS. Similarly, the high-frequency IP and additional components (HFC1 and HFC2) are most likely unrelated to the low-frequency IP and originate from different regions and physical processes. All these relatively recent findings demonstrate that, despite being one of the longest studied pulsars, Crab pulsar has still many mysteries, which can teach us a lot about the entire pulsar population. We note, however, that the micro and nano-second structure of Crab GPs cannot be observed at the low frequencies of the presented study due to the temporal scatter broadening caused by the ISM and Crab Nebula.
@@
After the discovery of fast radio bursts (2007Sci…318..777L, 52, FRBs;), GPs from extra-galactic pulsars were hypothesised to be responsible for at least some types (i.e. repeating) of FRBs 2016MNRAS.457..232C (17, 58), and Crab GPs have been extensively studied in the context of FRBs (e.g. 2022NatAs…6..393N (66, 35)) as well as for testing this hypothesis (e.g. 2024ApJ…973…87D (21, 5, 63)). Based on the energy distribution of Crab GPs and other characteristics (e.g. time between pulses), these studies have disfavoured the theory that all FRBs are caused by GPs from extra-galactic pulsars, and it is more likely that majority of FRBs are caused by objects like magnetars 2019MNRAS.485.4091M (62, 27). However, some weaker FRBs, such as FRB 20200120E, could be caused by extra-galactic Crab-like pulsars (see 2022NatAs…6..393N (66)). For the most recent review of the FRB field see 2022A&ARv..30….2P (71), while FRB models are collated at https://frbtheorycat.org/.
@@
While the discovery and early Crab GP studies (e.g. 1995ApJ…453..433L (54)) were performed at frequencies below 1 GHz, the majority of later investigations were conducted at frequencies above 1 GHz (e.g. 2024ApJ…973…87D (21, 5, 22, 10)) up to even 15.1 GHz (e.g. 2010A&A…524A..60J (37, 32)). On the high end of electromagnetic spectrum GP studies extend to -rays; see 2021Univ….7..448A (2) for the most recent review, and 2021Sci…373..425L (49, 1) for recent results at PeV energies.
@@
The advent of new (LOFAR 2013A&A…556A…2V (87), MWA 2013PASA…30….7T (86, 90), LWA 2013ITAP…61.2540E (25) etc.) and upcoming instruments, like low-frequency Square Kilometre Array telescope (2009IEEEP..97.1482D, 19, SKA-Low;)111https://www.skatelescope.org/ with large fields-of-view (FoVs), has begun to unlock the wealth of information available at these frequencies, and reinvigorated pulsar studies at frequencies below 400 MHz. Several recent GP studies were performed down to 110 MHz (e.g. 2012A&A…538A…7K (39, 63, 68, 64, 11)). They addressed various aspects of Crab GPs ranging from monitoring plasma in the surrounding environment and along the line of sight (LoS) to the pulsar, through variations in dispersion measure (DM) and scatter broadening time to statistical properties of GPs, their emission mechanisms and similarities to FRBs.
@@
This paper presents results from one such instrument, the Engineering Development Array version 2 (2022JATIS…8a1010W, 89, EDA2;) a prototype station of the SKA-Low. Despite modest sensitivity we were able to record the largest ever sample (95000) of Crab giant pulses observed at SKA-Low frequencies, superseding even the largest ever sample of 30000 Crab GPs recorded by 1995ApJ…453..433L (54). The EDA2 was originally deployed as a risk mitigation and reference system to the official SKA-Low prototype station the Aperture Array Verification System 2 station (2022JATIS…8a1014M, 59, AAVS2;). Real-time station beamforming and pulsar observing mode were commissioned at an early stage and already resulted in early SKA-Low pulsar science results 2022PASA…39…42L (48, 47). Both stations have also been used for all-sky transient searches 2021PASA…38…23S (80).
@@
Recently, a new pipeline, based on a code developed for the MWA by 2024PASA…41…11S (78), has been implemented and deployed for non-targeted (“blind”) searches and monitoring of repeating FRBs, magnetars, RRATs and pulsars with a single station beam of EDA2. In order to test and verify the pipeline, multiple observations of Crab pulsar were performed, and led to detection of thousands of GPs. These detections have confirmed that the station and the pipeline are capable of detecting highly dispersed pulses, and can be successful in searching for and/or monitoring FRBs, magnetars, RRATs or other fast transients. Especially, once the full FoV of the stations is unlocked by multi-beaming or high-time resolution all-sky imaging as proposed by 2022aapr.confE…1S (83).
@@ 2 EDA2 telescope and observations
The data for this study were recorded between 2024-12-14 and 2025-03-31 with the EDA2 station. The details of EDA2 can be found in the description paper 2022JATIS…8a1010W (89), and only a brief summary relevant to this project is provided here. The EDA2 has been a risk mitigation and a reference to the AAVS2 station, with the only major difference being the antenna design. The EDA2 comprises 256 MWA bow-tie dipoles, and was used as a reference to AAVS2 station (already de-commissioned) composed of 256 SKALA 4.1 antennas. Both instruments were located close to each other (some 500 m apart) at the SKA-Low site, Inyarrimanha Ilgari Bundara observatory222https://www.csiro.au/en/about/facilities-collections/mro. Analogue signals from individual antennas are digitised and the entire bandwidth (400 MHz) channelised into 512 coarse channels in the backend Tile Processing Units (2017JAI…..641014N, 65, TPMs;). One of the main tasks of TPMs is to form real-time station beams in 1.08 usec time resolution over the entire SKA-Low observing bandwidth (50 – 350 MHz).
@@ 3 Data processing pipeline
The recorded DADA files have been processed with eda2frb pipeline333https://github.com/marcinsokolowski/eda2frb developed for this project (Fig. 1). The pipeline has been based on the FRB search pipeline applied to MWA incoherent beam 2024PASA…41…11S (78)444athttps://github.com/marcinsokolowski/mwafrb. It has been executed on a server computer installed at the observatory and consists of the following main steps. First, complex voltages in each coarse channel DADA file are fine channelised into 64 fine channels (approximately 14.5 kHz each) by a custom-developed spectrometer using FFTW (CPU) or cuFFT (GPU) libraries (the choice is optional). The resulting spectra (in 69.12 usec time resolution) are averaged to 0.96768 ms time resolution, and saved to separate filterbank files. Then, 40 filterbank files from each coarse channel are merged into a single filterbank covering the entire observing bandwidth (31.25 MHz). These steps are represented by blue boxes in Figure 1.
The single pulse search is performed on the resulting wideband filterbank file using both PRESTO 2011ascl.soft07017R (73) and FREDDA 2019ascl.soft06003B (4) software packages. One of the PRESTO executions was performed at Crab pulsar’s DM=56.72 pc/cm3 as measured during this period by the Jodrell Bank observatory555https://www.jb.man.ac.uk/pulsar/crab/crab2.txt. Selected FREDDA candidates were visually inspected, and an example of a GP with the highest signal-to-noise (SNR) ratio of 354 recorded in this campaign is shown in Figure 2. This is the brightest pulse in the sample with flux density of 16.6 kJy and fluence 76 Jy ms. We note that the pulse broadening effects were not removed in our analysis (e.g. via de-convolution, 2003ApJ…584..782B (7)), which may lead to a factor of 2 underestimation of the intrinsic peak flux density (see discussion in 2007ApJ…665..618B (11)).
It is worth mentioning that some previous studies recorded low-frequency (600 MHz) GPs upon receiving a trigger from high-frequency (1400 MHz) detections (e.g. 1999ApJ…517..460S (77)). However, in our study observations were performed at a single frequency (31.25 MHz bandwidth centred on 215 MHz), which was not strongly impacted by temporal pulse broadening, ionospheric scintillation or RFI 2016ApJ…829…62E (23). The moderate data rates ( 0.5 TB/h) enabled us to record 1 – 2 h observations. Hence, we were able to identify GPs independently (without relying on any external information) by performing a single pulse search offline using the standard software packages (i.e. PRESTO and FREDDA).
@@ 4.1 Post-processing of PRESTO output
Merging candidates - a single GP may result in multiple candidates with different arrival times, DMs and SNRs. Therefore, in order to avoid double-counting, the candidates were merged into a single candidate with maximum SNR within a 45 ms time window as highly scattering-broadened pulses can extend up to this duration. In principle, we could have missed two GPs within one pulsar period (33 ms), but the results of merging were visually inspected and such cases have never been found. In fact, so far, the so-called double (or “twin”) GPs, from MP and IP within the same pulsar period, have only been reported in L-band 2010A&A…515A..36K (40). Also, recently, 2023ApJ…959..111L (50) reported detections of clusters of micro-bursts (even up to 6) within a single pulse window (1 ms), but such instances cannot be identified in our data due to scatter broadening at our frequencies and 1 ms time resolution.
In order to find double-GPs, we performed a dedicated search in every recorded de-dispersed time series, and did not find any instance of two pulses separated by less than a pulsar period (33 ms). Hence, we are confident that the merging procedure does not affect the statistical analysis of fluence and flux density distributions, or any other parts of the presented analysis.